OpenStudy (goldendragon01):

Fan and Medal if right. Which equations in point-slope form are equations of the line that passes through the point (4, 5) and (3, 1)? Choose exactly two answers that are correct. A. y-3=7/6(x-1) B. y-5=6/7(x-4) C. y+1=6/7(x+3) D. y-4=7/6(x-5)

OpenStudy (shinalcantara):

getting the slope of the line you'll have m = (y-y1)/(x-x1) get the slope using the formula. just substitute (4, 5) and (3,1)

OpenStudy (hauntedwoodsgal):

wow yeh your smart

OpenStudy (shinalcantara):

after getting the slope, all you have to do is substitute it to the slope form (y-y1) = m(x-x1) where m is the slope. you can use either (4,5) or (3,1) for (x1,y1) in the substitution

OpenStudy (goldendragon01):

ok ty